# Norming Markushevich bases: recent results and open problems

### v1

@@ 3. Classes of non-separable Banach spaces and M-bases
It is clear that a functional $\varphi\in\mathcal{X}^{*}$ is countably supported if and only if $\mathsf{ev}(\varphi)\in\Sigma(\Gamma)$. Hence if a Banach space $\mathcal{X}$ is $1$-Plichko, the set of countably supported functionals in $B_{\mathcal{X}^{*}}$ is $w^{*}$-dense in the dual ball by (2.1) and the map $\mathsf{ev}$ witnesses that $(B_{\mathcal{X}^{*}},w^{*})$ is Valdivia. However, it is not true that the dual ball of $\mathcal{X}$ is Valdivia when $\mathcal{X}$ is merely Plichko, [60], see Remark 5.3. Further, it is also not true that $\mathcal{X}$ must be $1$-Plichko if $(B_{\mathcal{X}^{*}},w^{*})$ is Valdivia, [63]. It seems to be still open if $\mathcal{X}$ must be Plichko under this assumption, [63]. Finally, when $\mathcal{X}$ is a $\mathcal{C}(\mathcal{K})$ space, $\mathcal{C}(\mathcal{K})$ is $1$-Plichko for $\mathcal{K}$ Valdivia, [80] (see [61, Proposition 5.1]), while the converse fails, [10], [64, Section 4]. Moreover, if $\mathcal{K}$ has a dense set of $G_{\delta}$ points, $\mathcal{K}$ is Valdivia if and only if $\mathcal{C}(\mathcal{K})$ is $1$-Plichko, if and only if $(B_{\mathcal{C}(\mathcal{K})^{*}},w^{*})$ is Valdivia, [61, Theorem 5.3].
@@
To summarise, we have the following implications, no one of which can be reversed:
@@
In the last part of this section, we briefly mention some properties of Asplund spaces. The class of Asplund spaces is transversal to the previous ones and it does not constitute a generalisation of separable Banach spaces; yet, it seems most appropriate to recall it here.
The original definition is that a Banach space $\mathcal{X}$ is Asplund if every convex continuous function defined on a convex open subset $\mathcal{U}$ of $\mathcal{X}$ is Fréchet differentiable on a dense $G_{\delta}$ subset of $\mathcal{U}$. However, it is nowadays customary to define a Banach space $\mathcal{X}$ to be Asplund if every separable subspace of $\mathcal{X}$ has a separable dual. This equivalent definition has the advantage to be the shortest to formulate and to readily give several properties and examples. For instance, it readily implies that Asplund spaces are closed under taking closed subspaces and quotients. Moreover, it shows that every reflexive Banach space, as well as $c_{0}(\Gamma)$ for all set $\Gamma$, is an example of an Asplund space; on the other hand, $\ell_{1}(\Gamma)$ is an easy example of Banach space that is not Asplund. The equivalence between the above definitions can be found for example in [23, Theorem I.5.7].
When considering $\mathcal{C}(\mathcal{K})$ spaces, a $\mathcal{C}(\mathcal{K})$ space is Asplund if and only if the compact $\mathcal{K}$ is scattered, see, e.g., [23, Lemma VI.8.3]. Recall that a compact space $\mathcal{K}$ is scattered if every closed subset of $\mathcal{K}$ admits an isolated point. In particular, a very important example for us is that $\mathcal{C}([0,\omega_{1}])$ is an Asplund space. For further information and historical background on Asplund spaces we refer to [23, § 1.5], [26], [40], [81], and the references therein.
@@ 4. Norming M-bases and WCG spaces
In order to explain the connection between WCG Banach spaces and norming M-bases, we need two notions, that of projectional resolution of the identity (PRI for short) and that of locally uniformly rotund (LUR, for short) norm. A PRI in a Banach space $\mathcal{X}$ is a well-ordered sequence of norm-one projections on $\mathcal{X}$ onto smaller subspaces, that are moreover compatible and that satisfy a certain continuity condition. For the precise definition we refer the reader to [23, Section VI.1], or [40, Section 6.1.4]. For us it will be more useful to keep in mind the idea that a PRI is way to decompose a Banach space into smaller well-behaved pieces, from which information on the space can be reconstructed, arguing by transfinite induction on the density character of the space. PRI’s were first constructed by Lindenstrauss [71] in every reflexive Banach space and shortly after by Amir and Lindenstrauss in every WCG Banach space [5]. Since then, their use in non-separable Banach spaces has been pervasive and their usefulness cannot be overstated. Shortly after their introduction, Troyanski obtained his celebrated renorming theorem [96] that every WCG Banach space admits an LUR norm. Let us just say here that, besides being among the most important notions in renorming, LUR norms are also extremely useful outside the isomorphic theory; for instance they are used in Kadets’ proof that all separable Banach spaces are homeomorphic, see e.g., [40, Section 9.5.3].
Subsequently, John and Zizler [53] constructed a PRI in every Banach space with $1$-norming M-basis and deduced that Banach spaces with norming M-basis admit an LUR renorming. It was then natural to conjecture a connection between WCG spaces and norming M-bases; the most optimistic hope that a Banach space is WCG if and only if it admits a norming M-basis is however too naive due to the example of $\ell_{1}(\Gamma)$. The validity of the converse implication was asked in the same paper [53] and eventually became one of the classical problems in Banach space theory that was reiterated multiple times in several articles and books, [39, Problem 111], [43, p. 173], [62], [83, p. 108], or [103, Problem 10].
@@
Notice that in Problem 4.1 it is sufficient to consider $\mathcal{C}(\mathcal{K})$ spaces. Indeed, if $\mathcal{X}$ is any WCG Banach space, then $(B_{\mathcal{X}^{*}},w^{*})$ is Eberlein, hence $\mathcal{C}(B_{\mathcal{X}^{*}})$ is WCG as well, as we saw in Section 3. Moreover, $\mathcal{X}$ is a subspace of $\mathcal{C}(B_{\mathcal{X}^{*}})$. As we will see in Theorem 6.3 below, a WCG subspace of a Banach space with $\lambda$-norming M-basis admits such an M-basis as well. Thus a positive answer to the problem for $\mathcal{C}(\mathcal{K})$ spaces implies a positive answer for all WCG Banach spaces.
@@
As we will see, the role of the Asplund property essentially is to obtain information on the dual space as well. To explain this, let us start with a result characterising spaces with a shrinking M-basis (the precise attribution to the result will be given below).
@@
We now discuss the contributions that eventually led to the characterisation in Theorem 4.3. To begin with, the implications (v)$\implies$(vi)$\implies$(iv) are standard [23] (and don’t require the WLD assumption), while (ii)$\implies$(iii)$\implies$(iv) is obvious. Next, (i)$\implies$(ii) is elementary and we shall explain it here (see also the remarks after [26, Definition 6.2.3]). If $\{e_{\alpha};\varphi_{\alpha}\}_{\alpha\in\Gamma}$ is a shrinking M-basis for $\mathcal{X}$ and $\|e_{\alpha}\|=1$, then $\{e_{\alpha}\}_{\alpha\in\Gamma}\cup\{0\}$ is weakly compact: in fact, to check that every subsequence of $\{e_{\alpha}\}_{\alpha\in\Gamma}$ converges weakly to 00 it is enough to test on the linearly dense set $\{\varphi_{\alpha}\}_{\alpha\in\Gamma}$, for which functionals the convergence is obvious. Hence, $\mathcal{X}$ is WCG. Next, if $\mathcal{Z}$ is a separable subspace of $\mathcal{X}$, take a countable set $\Gamma_{0}\subseteq\Gamma$ with $\mathcal{Z}\subseteq\overline{\rm span}\{e_{\alpha}\}_{\alpha\in\Gamma_{0}}$; then $\mathcal{Z}^{*}\subseteq\overline{\rm span}\{\varphi_{\alpha}\mathord{\upharpoonright}_{\mathcal{Z}}\}_{\alpha\in\Gamma_{0}}$. In fact, take a functional $\varphi\in\mathcal{Z}^{*}$, extend it to a functional on $\mathcal{X}$, and approximate the extension by a linear combination $\sum_{\alpha\in\Gamma}c_{\alpha}\varphi_{\alpha}$; then $\sum_{\alpha\in\Gamma}c_{\alpha}\varphi_{\alpha}\mathord{\upharpoonright}_{\mathcal{Z}}\in{\rm span}\{\varphi_{\alpha}\mathord{\upharpoonright}_{\mathcal{Z}}\}_{\alpha\in\Gamma_{0}}$ approximates $\varphi$ (note that $\varphi_{\alpha}\mathord{\upharpoonright}_{\mathcal{Z}}=0$ for $\alpha\notin\Gamma_{0}$).
@@
Now, to the essential contributions. Fabian [25] proved that a WCD Asplund space is actually WCG and its dual admits a dual LUR norm. Inspection of the proof shows that Fabian built a shrinking M-basis and derived the WCG property out of it. Hence, (i), (ii), and (iii) are equivalent; moreover, they are also equivalent to (iv), (v), and (vi) if one replaces WLD with WCD in them. The last result needed to close the circle of implications is that WLD Asplund spaces admit a shrinking M-basis, that was proved by Valdivia [97, Proposition 2]. Let us also point out the earlier result due to John and Zizler [54, Theorem 1] that $\mathcal{X}$ has a shrinking M-basis if and only if $\mathcal{X}$ is WCG and admits a Fréchet smooth norm, if and only if $\mathcal{X}$ is WCG and $\mathcal{X}^{*}$ admits a dual LUR norm. We refer to [43, Theorem 6.3], or [26, Theorem 8.3.3] for a proof of Theorem 4.3.
@@
Next, we come to the role of Theorem 4.3 for norming M-bases. Fabian’s argument consists in working in $\mathcal{X}$ and in $\mathcal{X}^{*}$ simultaneously to build a PRI in $\mathcal{X}$ such that the adjoints of the projections form a PRI in $\mathcal{X}^{*}$ (let us call a PRI with this property a shrinking PRI). Then this combines with a result of Vašák [101, Theorem 2] who built a shrinking M-basis from a shrinking PRI. Similarly, Valdivia built a shrinking PRI in WLD Asplund spaces, [97, Lemma 2]. Moreover, at the same time of [25], Fabian and Godefroy [27] built a PRI in the dual of every Asplund space. Therefore, the Asplund assumption gives a PRI in the dual, while the WCD one gives a PRI in the space. Fabian’s method consisted in merging the two techniques to obtain compatible projections in the space and in the dual simultaneously (see the explanation in [26, Chapter 8] or [23, Sections VI.2–VI.4]). However, John and Zizler [53] built a PRI in presence of a $1$-norming M-basis in $\mathcal{X}$. Therefore, it was natural to conjecture that one could also glue the techniques from John–Zizler [53] with these from Fabian–Godefroy [27] to build a shrinking PRI in Asplund spaces with a norming M-basis. That it were the case was indeed conjectured by Godefroy at the times when [23] was in preparation and the problem was subsequently recorded in various articles and books, see, e.g., [3], [43, p. 211], [39, Problem 112].
@@
There exists a uniform Eberlein compactum $\mathcal{K}$ such that the Banach space $\mathcal{C}(\mathcal{K})$ doesn’t admit any norming M-basis.
@@
We now explain some ingredients in the proof of Theorem 4.5. First of all, it is enough to build a uniform Eberlein compact space $\mathcal{K}_{n}\subseteq\ell_{2}(\omega_{1})$ such that $\mathcal{C}(\mathcal{K}_{n})$ doesn’t admit a $1/n$-norming M-basis. Indeed, up to a scaling, one can assume that $\mathcal{K}_{n}\subseteq\frac{1}{n}B_{\ell_{2}(\{n\}\times\omega_{1})}\subseteq\ell_{2}(\mathbb{N}\times\omega_{1})$ and that $\mathcal{K}_{n}$ contains the origin. Then
@@
For every $n\in\mathbb{N}$ there exists a uniform Eberlein compactum $\mathcal{K}_{n}\subseteq\ell_{2}(\omega_{1})$ such that $\mathcal{C}(\mathcal{K}_{n})$ doesn’t admit a $1/n$-norming M-basis.
@@
such that $t_{j}\in D_{j}$ for all $j$ and such that the sets $H(\alpha_{1},\dots,\alpha_{k})$ are mutually disjoint. This is easily done by transfinite induction: once all sets $H(\alpha_{1},\dots,\alpha_{k})$ have been chosen for all $\alpha_{1},\dots,\alpha_{k}\leqslant\beta$, only countably many points have been taken in each $D_{j}$, that has cardinality continuum. Hence there is room for the next choice.
@@
We now pass to the negative solution to Problem 4.4. The main part of the argument consists in building a certain compact topological space $\mathcal{K}_{\varrho}$, that we will also mention in Section 7; then the desired counterexample will be a subspace of $\mathcal{C}(\mathcal{K}_{\varrho})$.
@@
It is also convenient to add the ‘boundary condition’ $\varrho(\alpha,\alpha)=0$. Conditions ($\varrho$1) and ($\varrho$2) above imply that $\varrho$ resembles in certain ways a metric on $\omega_{1}$, whence the name ordinal metric. They were introduced by Todorčević in [94] and analysed in detail in [95]; see also [11]. The construction of a function $\varrho$ as in Proposition 4.8 can be found in [95, Lemma 3.2.2] and some details are also in [45, Section 3].
@@
Notice that $\mathcal{K}_{\varrho}$ is the closure of a collection of finite sets, hence elements of $\mathcal{K}_{\varrho}$ can be approximated by finite sets; in this sense, $\mathcal{K}_{\varrho}$ resembles the Tikhonov cube $[0,1]^{\omega_{1}}$, or the Cantor cube $\{0,1\}^{\omega_{1}}$. On the other hand, $\mathcal{K}_{\varrho}$ contains very few infinite sets, just the initial intervals, for which reason it is scattered. In this respect, it resembles the behaviour of $[0,\omega_{1}]$ (we will return on this point in Remark 7.2).
@@
The validity of (i) and (ii) follows easily from what has been said before, while (iv) can be deduced from (iii) via a certain maximality argument, somewhat similar to the proof that Eberlein subsets of $\{0,1\}^{\Gamma}$ are scattered (as in [43, Lemma 2.53]).
@@
Here, the main point consists in proving the crucial formula
@@
Therefore, we are only left with the proof of (4.1). For this, it is clearly enough to check that the two functionals agree on each $f_{\beta}$, $\beta<\omega_{1}$. However, this is immediate: if $\beta\in A$, both functionals equal $1$, if $\beta\notin A$ both vanish. ∎
@@
In conclusion to this section, let us point out explicitly that the counterexample in Theorem 4.10 is not a $\mathcal{C}(\mathcal{K})$ space, unlikely the example of Theorem 4.5. In particular, we don’t know the answer to the following problem.

### v2

@@ 3. Classes of non-separable Banach spaces and M-bases
It is clear that a functional $\varphi\in\mathcal{X}^{*}$ is countably supported if and only if $\mathsf{ev}(\varphi)\in\Sigma(\Gamma)$. Hence if a Banach space $\mathcal{X}$ is $1$-Plichko, the set of countably supported functionals in $B_{\mathcal{X}^{*}}$ is $w^{*}$-dense in the dual ball by (2.1) and the map $\mathsf{ev}$ witnesses that $(B_{\mathcal{X}^{*}},w^{*})$ is Valdivia. However, it is not true that the dual ball of $\mathcal{X}$ is Valdivia when $\mathcal{X}$ is merely Plichko, [67], see Remark 5.3. Further, it is also not true that $\mathcal{X}$ must be $1$-Plichko if $(B_{\mathcal{X}^{*}},w^{*})$ is Valdivia, [72]. It seems to be still open if $\mathcal{X}$ must be Plichko under this assumption, [72]. Finally, when $\mathcal{X}$ is a $\mathcal{C}(\mathcal{K})$ space, $\mathcal{C}(\mathcal{K})$ is $1$-Plichko for $\mathcal{K}$ Valdivia, [91] (see [70, Proposition 5.1]), while the converse fails, [10], [73, Section 4]. Moreover, if $\mathcal{K}$ has a dense set of $G_{\delta}$ points, $\mathcal{K}$ is Valdivia if and only if $\mathcal{C}(\mathcal{K})$ is $1$-Plichko, if and only if $(B_{\mathcal{C}(\mathcal{K})^{*}},w^{*})$ is Valdivia, [70, Theorem 5.3].
@@
To summarise, we have the following implications, none of which can be reversed:
@@
In the last part of this section, we briefly mention some properties of Asplund spaces. The class of Asplund spaces is transversal to the previous ones and it does not constitute a generalisation of separable Banach spaces; yet, it seems most appropriate to recall it here. The original definition is that a Banach space $\mathcal{X}$ is Asplund if every convex continuous function defined on a convex open subset $\mathcal{U}$ of $\mathcal{X}$ is Fréchet differentiable on a dense $G_{\delta}$ subset of $\mathcal{U}$. However, it is nowadays customary to define a Banach space $\mathcal{X}$ to be Asplund if every separable subspace of $\mathcal{X}$ has a separable dual. This equivalent definition has the advantage to be the shortest to formulate and to directly give several properties and examples. For instance, it readily implies that Asplund spaces are closed under taking closed subspaces and quotients. Moreover, it shows that every reflexive Banach space, as well as $c_{0}(\Gamma)$ for all set $\Gamma$, is an example of an Asplund space; on the other hand, $\ell_{1}(\Gamma)$ is an easy example of Banach space that is not Asplund. The equivalence between the above definitions can be found for example in [28, Theorem I.5.7].
When considering $\mathcal{C}(\mathcal{K})$ spaces, a $\mathcal{C}(\mathcal{K})$ space is Asplund if and only if the compact $\mathcal{K}$ is scattered, see, e.g., [28, Lemma VI.8.3]. Recall that a compact space $\mathcal{K}$ is scattered if every closed subset of $\mathcal{K}$ admits an isolated point. In particular, a very important example for us is that $\mathcal{C}([0,\omega_{1}])$ is an Asplund space. For further information and historical background on Asplund spaces we refer to [28, § 1.5], [33], [47], [92], and the references therein.
@@ 4. Norming M-bases and WCG spaces
In order to explain the connection between WCG Banach spaces and norming M-bases, we need two notions, that of projectional resolution of the identity (PRI for short) and that of locally uniformly rotund (LUR, for short) norm. A PRI in a Banach space $\mathcal{X}$ is a well-ordered collection of norm-one projections on $\mathcal{X}$ onto smaller subspaces, that are moreover compatible and that satisfy a certain continuity condition. For the precise definition we refer the reader to [28, Section VI.1], or [47, Section 6.1.4]. For us it will be more useful to keep in mind the idea that a PRI is way to decompose a Banach space into smaller well-behaved pieces, from which information on the space can be reconstructed, arguing by transfinite induction on the density character of the space. PRI’s were first constructed by Lindenstrauss [81] in every reflexive Banach space and shortly after by Amir and Lindenstrauss in every WCG Banach space [5]. Since then, their use in non-separable Banach spaces has been pervasive and their usefulness cannot be overstated. Shortly after their introduction, Troyanski obtained his celebrated renorming theorem [107] that every WCG Banach space admits an LUR norm (see also Zizler’s extension [114]). Let us just say here that, besides being among the most important notions in renorming, LUR norms are also extremely useful outside the isomorphic theory; for instance they are used in Kadets’ proof that all separable Banach spaces are homeomorphic, see e.g., [47, Section 9.5.3].
Subsequently, John and Zizler [60] constructed a PRI in every Banach space with $1$-norming M-basis and deduced that Banach spaces with norming M-basis admit an LUR renorming. It was then natural to conjecture a connection between WCG spaces and norming M-bases; the most optimistic hope that a Banach space is WCG if and only if it admits a norming M-basis is however too naive due to the example of $\ell_{1}(\Gamma)$. The validity of the converse implication was asked in the same paper [60] and eventually became one of the classical problems in Banach space theory that was reiterated multiple times in several articles and books, [46, Problem 111], [50, p. 173], [71], [94, p. 108], or [115, Problem 10].
@@
Notice that in Problem 4.1 it is sufficient to consider $\mathcal{C}(\mathcal{K})$ spaces, in the sense that a positive answer for all WCG $\mathcal{C}(\mathcal{K})$ spaces would imply a positive answer for all WCG spaces. Indeed, if $\mathcal{X}$ is any WCG Banach space, then $(B_{\mathcal{X}^{*}},w^{*})$ is Eberlein, hence $\mathcal{C}(B_{\mathcal{X}^{*}})$ is WCG as well, as we saw in Section 3. Moreover, $\mathcal{X}$ is a subspace of $\mathcal{C}(B_{\mathcal{X}^{*}})$. As we will see in Theorem 6.3 below, a WCG subspace of a Banach space with $\lambda$-norming M-basis admits such an M-basis as well. Thus, the existence of a norming M-basis in $\mathcal{C}(B_{\mathcal{X}^{*}})$ would imply the existence of a norming M-basis in $\mathcal{X}$.
@@
As we will see, essentially the role of the Asplund property is to provide information on the dual space as well. To explain this, let us start with a result characterising spaces with a shrinking M-basis (the precise attribution to the result will be given below).
@@
We now discuss the contributions that eventually led to the characterisation in Theorem 4.3. To begin with, the implications (v)$\implies$(vi)$\implies$(iv) are standard [28] (and do not require the WLD assumption), while (ii)$\implies$(iii)$\implies$(iv) is obvious. Next, (i)$\implies$(ii) is elementary and we shall explain it here (see also the remarks after [33, Definition 6.2.3]). If $\{e_{\alpha};\varphi_{\alpha}\}_{\alpha\in\Gamma}$ is a shrinking M-basis for $\mathcal{X}$ and $\|e_{\alpha}\|=1$, then $\{e_{\alpha}\}_{\alpha\in\Gamma}\cup\{0\}$ is weakly compact: in fact, to check that every subsequence of $\{e_{\alpha}\}_{\alpha\in\Gamma}$ converges weakly to 00 it is enough to test on the linearly dense set $\{\varphi_{\alpha}\}_{\alpha\in\Gamma}$, for which functionals the convergence is obvious. Hence, $\mathcal{X}$ is WCG. Next, if $\mathcal{Z}$ is a separable subspace of $\mathcal{X}$, take a countable set $\Gamma_{0}\subseteq\Gamma$ with $\mathcal{Z}\subseteq\overline{\rm span}\{e_{\alpha}\}_{\alpha\in\Gamma_{0}}$; then $\mathcal{Z}^{*}\subseteq\overline{\rm span}\{\varphi_{\alpha}\mathord{\upharpoonright}_{\mathcal{Z}}\}_{\alpha\in\Gamma_{0}}$. In fact, take a functional $\varphi\in\mathcal{Z}^{*}$, extend it to a functional on $\mathcal{X}$, and approximate the extension by a linear combination $\sum_{\alpha\in\Gamma}c_{\alpha}\varphi_{\alpha}$; then $\sum_{\alpha\in\Gamma}c_{\alpha}\varphi_{\alpha}\mathord{\upharpoonright}_{\mathcal{Z}}\in{\rm span}\{\varphi_{\alpha}\mathord{\upharpoonright}_{\mathcal{Z}}\}_{\alpha\in\Gamma_{0}}$ approximates $\varphi$ (note that $\varphi_{\alpha}\mathord{\upharpoonright}_{\mathcal{Z}}=0$ for $\alpha\notin\Gamma_{0}$).
@@
Now, to the essential contributions. Fabian [31] proved that a WCD Asplund space is actually WCG and its dual admits a dual LUR norm. Inspection of the proof shows that Fabian built a shrinking M-basis and derived the WCG property out of it. Hence, (i), (ii), and (iii) are equivalent; moreover, they are also equivalent to (iv), (v), and (vi) if one replaces WLD with WCD in them. The last result needed to close the circle of implications is that WLD Asplund spaces admit a shrinking M-basis, that was proved by Valdivia [108, Proposition 2]. Let us also point out the earlier result due to John and Zizler [61, Theorem 1] that $\mathcal{X}$ has a shrinking M-basis if and only if $\mathcal{X}$ is WCG and admits a Fréchet smooth norm, if and only if $\mathcal{X}$ is WCG and $\mathcal{X}^{*}$ admits a dual LUR norm. We refer to [50, Theorem 6.3], or [33, Theorem 8.3.3] for a proof of Theorem 4.3.
@@
Next, we come to the role of Theorem 4.3 for norming M-bases. Fabian’s argument amounts to working in $\mathcal{X}$ and in $\mathcal{X}^{*}$ simultaneously to build a PRI in $\mathcal{X}$ such that the adjoints of the projections form a PRI in $\mathcal{X}^{*}$ (let us call a PRI with this property a shrinking PRI). Then this combines with a result of Vašák [112, Theorem 2] who built a shrinking M-basis from a shrinking PRI. Similarly, Valdivia built a shrinking PRI in WLD Asplund spaces, [108, Lemma 2]. Moreover, at the same time of [31], Fabian and Godefroy [34] built a PRI in the dual of every Asplund space. Therefore, the Asplund assumption gives a PRI in the dual, while the WCD one gives a PRI in the space. Fabian’s method consisted in merging the two techniques to obtain compatible projections in the space and in the dual simultaneously (see the explanation in [33, Chapter 8] or [28, Sections VI.2–VI.4]). However, John and Zizler [60] built a PRI in presence of a $1$-norming M-basis in $\mathcal{X}$. Therefore, it was natural to conjecture that one could also glue the techniques from John–Zizler [60] with these from Fabian–Godefroy [34] to build a shrinking PRI in Asplund spaces with a norming M-basis. That it were the case was indeed conjectured by Godefroy at the times when [28] was in preparation and the problem was subsequently recorded in various articles and books, see, e.g., [3], [50, p. 211], [46, Problem 112].
@@
There exists a uniform Eberlein compactum $\mathcal{K}$ such that the Banach space $\mathcal{C}(\mathcal{K})$ does not admit any norming M-basis.
@@
We now explain some ingredients in the proof of Theorem 4.5. First of all, it is enough to build a uniform Eberlein compact space $\mathcal{K}_{n}\subseteq\ell_{2}(\omega_{1})$ such that $\mathcal{C}(\mathcal{K}_{n})$ does not admit a $1/n$-norming M-basis. Indeed, up to a scaling, one can assume that $\mathcal{K}_{n}\subseteq\frac{1}{n}B_{\ell_{2}(\{n\}\times\omega_{1})}\subseteq\ell_{2}(\mathbb{N}\times\omega_{1})$ and that $\mathcal{K}_{n}$ contains the origin. Then
@@
For every $n\in\mathbb{N}$ there exists a uniform Eberlein compactum $\mathcal{K}_{n}\subseteq\ell_{2}(\omega_{1})$ such that $\mathcal{C}(\mathcal{K}_{n})$ does not admit a $1/n$-norming M-basis.
@@
with $t_{j}\in D_{j}$ for all $j$ and such that the sets $H(\alpha_{1},\dots,\alpha_{k})$ are mutually disjoint. This is easily done by transfinite induction: once all sets $H(\alpha_{1},\dots,\alpha_{k})$ have been chosen for all $\alpha_{1},\dots,\alpha_{k}\leqslant\beta$, only countably many points have been taken in each $D_{j}$, that has cardinality of the continuum. Hence there is room for the next choice.
@@
We now pass to the negative solution to Problem 4.4. The main part of the argument involves building a certain compact topological space $\mathcal{K}_{\varrho}$, that we will also mention in Sections 5 and 7; then the desired counterexample will be a subspace of $\mathcal{C}(\mathcal{K}_{\varrho})$.
@@
It is also convenient to add the ‘boundary condition’ $\varrho(\alpha,\alpha)=0$. Conditions ($\varrho$1) and ($\varrho$2) above imply that $\varrho$ resembles in certain ways a metric on $\omega_{1}$, whence the name ordinal metric. They were introduced by Todorčević in [105] and analysed in detail in [106]; see also [11]. The construction of a function $\varrho$ as in Proposition 4.8 can be found in [106, Lemma 3.2.2] and some details are also in [52, Section 3].
@@
Notice that $\mathcal{K}_{\varrho}$ is the closure of a collection of finite sets, hence elements of $\mathcal{K}_{\varrho}$ can be approximated by finite sets; in this respect, $\mathcal{K}_{\varrho}$ resembles the Tikhonov cube $[0,1]^{\omega_{1}}$, or the Cantor cube $\{0,1\}^{\omega_{1}}$. On the other hand, $\mathcal{K}_{\varrho}$ contains very few infinite sets, just the initial intervals, for which reason it is scattered. This property resembles the behaviour of $[0,\omega_{1}]$ (we will return to this point in Remark 7.2).
@@
The validity of (i) and (ii) follows easily from what has been said before, while (iv) can be deduced from (iii) via a certain maximality argument, somewhat similar to the proof that Eberlein subsets of $\{0,1\}^{\Gamma}$ are scattered (as in [50, Lemma 2.53]).
@@
Here the main point lies in proving the crucial formula
@@
Therefore, we are only left with the proof of (4.1). For this, it is clearly enough to check that the two functionals agree on each $f_{\beta}$, $\beta<\omega_{1}$. However, this is immediate: if $\beta\in A$, both functionals equal $1$, if $\beta\notin A$, both vanish. ∎
@@
To conclude our discussion of [52], let us point out explicitly that the counterexample in Theorem 4.10 is not a $\mathcal{C}(\mathcal{K})$ space, unlikely the example of Theorem 4.5. In particular, we do not know the answer to the following problem.